OpenStudy (anonymous):

What is the solution set for |–4x – 8| = 4? What is the solution set for |–4x – 8| = 4? @Mathematics

OpenStudy (diyadiya):

-4x-8=4 -4x=12 x=-3 -(-4x-8)=4 4x+8=4 4x=-4 x=-1 x=-3 x=-1